OpenStudy (anonymous):

Find the volume of a cone with a base radius of 4yd and a height of 12yd . Write the exact volume in terms of π , and be sure to include the correct unit in your answer.

OpenStudy (jhannybean):

Volume of a cone:\[V=\pi r^2 \left(\frac{h}{3}\right)\]

OpenStudy (anonymous):

201.06

OpenStudy (anonymous):

is it cubed or squared

OpenStudy (jhannybean):

So \[V= \pi (4yd)^2\left(\frac{12yd}{3}\right)\]

OpenStudy (jhannybean):

And I don't know the answer :P Just working off from what you have in your equation so I can't really confirm your answer, haha.

hartnn (hartnn):

yd is yards you can ignore that while using the equation \(V= \pi (4)^2\left(\frac{12}{3}\right)\)

OpenStudy (jhannybean):

Oh was that yards

OpenStudy (jhannybean):

Haha whoops xD

OpenStudy (anonymous):

I plugged in the numbers and got 201.06 but for my answer i have to have the correct unit

hartnn (hartnn):

don't put the value of pi

hartnn (hartnn):

just evaluate \(\Large 4^2 \times \dfrac{12}{3}\)

OpenStudy (jhannybean):

You're solving in terms of pi, meaning youll have some number * \(\pi\)

OpenStudy (jhannybean):

as your answer.

OpenStudy (anonymous):

oh so i take 4^2 multiplied by 12/3 then

hartnn (hartnn):

what did u get ?

OpenStudy (anonymous):

64

hartnn (hartnn):

correct! so your volume is \(64 \pi~~ yd^3 \)

OpenStudy (anonymous):

Thank you!

hartnn (hartnn):

welcome ^_^